OpenStudy (anonymous):

the taxi fare in a city is; for the first km=rs 50 and for subsequent distance=rs20 taking the total distance covered as 2 km and the total fare as rs y. write a linear equation

OpenStudy (tyteen4a03):

Suppose rs is a currency name, the equation is \(50 + 20x = y\) where the subsequent distance is x km. Plug in the subsequent value of 1km (first km is already charged) into the equation and find the answer.

terenzreignz (terenzreignz):

Alternatively, it just be y = 50 + 20(x - 1) That way, entering 1km also results in the actual fare.

OpenStudy (anonymous):

but the distance is given = 2km

terenzreignz (terenzreignz):

Yeah, so input 2km y = 50 + 20(2 - 1) y = 50 + 20(1) y = 70 And that's the fare for 2 km

OpenStudy (anonymous):

yup...that i m getting too. same asnwer